# Computing Minimal Absent Words and Extended Bispecial Factors with CDAWG Space

### v1

@@ Abstract
A string $w$ is said to be a minimal absent word (MAW) for a string $S$ if $w$ does not occur in $S$ and any proper substring of $w$ occurs in $S$. We focus on non-trivial MAWs which are of length at least 2. Finding such non-trivial MAWs for a given string is motivated for applications in bioinformatics and data compression. Fujishige et al. [TCS 2023] proposed a data structure of size $\Theta(n)$ that can output the set $\mathsf{MAW}(S)$ of all MAWs for a given string $S$ of length $n$ in $O(n+|\mathsf{MAW}(S)|)$ time, based on the directed acyclic word graph (DAWG). In this paper, we present a more space efficient data structure based on the compact DAWG (CDAWG), which can output $\mathsf{MAW}(S)$ in $O(|\mathsf{MAW}(S)|)$ time with $O(\mathsf{e}_{\min})$ space, where $\mathsf{e}_{\min}$ denotes the minimum of the sizes of the CDAWGs for $S$ and for its reversal $S^{R}$. For any strings of length $n$, it holds that $\mathsf{e}_{\min}<2n$, and for highly repetitive strings $\mathsf{e}_{\min}$ can be sublinear (up to logarithmic) in $n$. We also show that MAWs and their generalization minimal rare words have close relationships with extended bispecial factors, via the CDAWG.
@@ 1 Introduction
A string $w$ is said to be a minimal absent word (MAW) for a string $S$ if $w$ does not occur in $S$ and any proper substring of $w$ occurs in $S$. Throughout this paper, we focus on MAWs which are of form $aub$, where $a,b$ are characters and $u$ is a (possibly empty) string. Finding such MAWs for a given string is motivated for applications including bioinformatics [1, 12, 23, 19] and data compression [14, 15, 3].
@@
In this paper, we present a more space efficient data structure based on the compact DAWG (CDAWG[10], which takes $O(\mathsf{e}_{\min})$ space and can output $\mathsf{MAW}(S)$ in $O(|\mathsf{MAW}(S)|)$ time, where $\mathsf{e}_{\min}=\{\mathsf{el}(S),\mathsf{er}(S)\}$, with $\mathsf{el}(S)$ and $\mathsf{er}(S)$ being the numbers of edges of the CDAWGs for $S$ and for its reversal $S^{R}$, respectively. For any string $S$ of length $n$ it holds that the number of edges of the CDAWG for $S$ is less than $2n$ and it can be sublinear (up to logarithmic) in $n$ for highly repetitive strings [25, 24, 5].
@@ 2.2 MAW, MUS, MRW, and EBF
A string $w=aub$ with $a,b\in\Sigma$ and $u\in\Sigma^{*}$ is called an extended bispecial factor (EBF) for another string $S$ if $\mathsf{occ}_{S}(a^{\prime}u)\geq 1$ and $\mathsf{occ}_{S}(ub^{\prime})\geq 1$ for some characters $a^{\prime}\neq a$ and $b^{\prime}\neq b$. Let $\mathsf{EBF}(S)$ denote the set of EBFs for string $S$.
@@ 4.1 Computing MAWs
Let $u$ be any node in $\mathsf{CDAWG}(S)$ and suppose it has $k$ in-coming edges. The above process of adding secondary edges to $\mathsf{LPT}(S)$ can be regarded as splitting $u$ into $k+1$ nodes $u^{\prime},\ell_{1},\ldots,\ell_{k}$, such that $|\mathsf{str}(u^{\prime})|<|\mathsf{str}(\ell_{1})|<\cdots<|\mathsf{str}(\ell_{k})|$. We define the suffix links of these nodes such that $\mathsf{slink}(u^{\prime})=\ell_{1}$, $\mathsf{slink}(\ell_{i-1})=\ell_{i}$ for $1\leq i<k$, and $\mathsf{slink}(\ell_{k})=v^{\prime}$ such that $v^{\prime}$ is the node in $\mathsf{LPT}^{+}(S)$ that corresponds to the destination node $v=\mathsf{slink}(u)$ of the suffix link of $u$ in $\mathsf{CDAWG}(S)$. We call these copied nodes $\ell_{1},\ldots,\ell_{k}$ as gray nodes, and the other node $u^{\prime}$ as a white node. For convenience, we assume that the source is a white node. See Fig. 3 for illustration.
@@
For $i=1$, node $u_{1}~{}(=u)$ has a hard Weiner link to $\hat{u}$. Let $a$ be the character label of this hard Weiner link. Let $\mathsf{child}(\hat{u})$ and $\mathsf{child}(\mathsf{slink}(u))=\mathsf{child}(u_{1})$ be the sorted lists of the first characters of the out-edge labels from $\hat{u}$ and from $u_{1}$. Note that $\mathsf{child}(\hat{u})\subseteq\mathsf{child}(u_{1})$. Then, it follows from the definition of MAWs and Lemma 2 that $au_{1}b\in\mathsf{MAW}(S)$ iff $b\in\mathsf{child}(u_{1})\setminus\mathsf{child}(\hat{u})$.
For each $i=2,\ldots,m$, observe that node $u_{i}$ has a (virtual) soft Weiner link with the same character label $a$, and leading to the same node $\hat{u}$. We emphasize that this soft Weiner link only virtual and is not stored in our $O(\mathsf{er}(S))$-space data structure. We have the following lemma:
@@
Let $c$ be the first character of the label of the edge $(u_{i+1},u_{i})$. Let $\hat{u}_{i}$ denote the (virtual) implicit node on the edge $(v,u)$, such that $\mathsf{slink}(\hat{u}_{i})=u_{i}$. Note that $c$ is the unique character that immediately follows the locus of $\hat{u}_{i}$ in the string label of edge $(v,u)$. Since $u_{i}$ is branching but $\hat{u}_{i}$ is non-branching, the set $\mathsf{child}(u_{i})\setminus\mathsf{child}(\hat{u}_{i})=\mathsf{child}(u_{i})\setminus\{c\}$ is not empty. Thus, for any character $b\in\mathsf{child}(u_{i})\setminus\{c\}$, we have $au_{i}b\in\mathsf{MAW}(S)$. ∎
@@
Let us analyze the time complexity and the data structure space usage. In the case with $i\geq 2$, by Lemma 4, accessing each node $u_{i}$ can be charged to any one of the MAWs reported with $u_{i}$. The time required for $u_{i}$ is thus linear in the number of reported MAWs $au_{i}b$. In the case with $i=1$, we compare the sorted lists $\mathsf{child}(\hat{u})$ and $\mathsf{child}(u_{1})$. Each successful character comparison with a character $b\in\mathsf{child}(u_{1})\setminus\mathsf{child}(\hat{u})$ for which MAW $au_{1}b$ is reported. On the other hand, each unsuccessful character comparison with a character $b\in\mathsf{child}(u_{1})\cap\mathsf{child}(\hat{u})$ can be charged to the corresponding out-edge of $\hat{u}$ whose string label begins with $b$. Thus, we can output all MAWs in $O(\mathsf{er}(S)+|\mathsf{MAW}(S)|)$ time. Further, when $|\mathsf{MAW}(S)|\leq\mathsf{er}(S)$, then we can store the set $\mathsf{MAW}(S)$ with $O(|\mathsf{MAW}(S)|)\subseteq O(\mathsf{er}(S))$ space and output each MAW in $O(1)$ time, using Lemma 3. This leads to our $O(|\mathsf{MAW}(S)|)$-time $O(\mathsf{er}(S))$-space data structure.
@@
For each node $u_{i}$ of $\mathsf{LPT}^{+}(S)$, the number of distinct characters $a$ such that $au_{i}b\in\mathsf{MAW}(S)$ is at most the number of (hard or sort) Weiner links from $u_{i}$, which is bounded by $\sigma$. The number of distinct characters $b$ such that $au_{i}b\in\mathsf{MAW}(S)$ does not exceed the number of out-edges from $u_{i}$. By summing them up for all nodes $u_{i}$ of $\mathsf{LPT}^{+}(S)$, we obtain $|\mathsf{MAW}(S)|=O(\sigma\cdot\mathsf{er}(S))$. Since the same argument holds for the reversed string $S^{R}$ and $\mathsf{LPT}^{+}(S^{R})$, we have $|\mathsf{MAW}(S)|=O(\sigma\cdot\min\{\mathsf{er}(S),\mathsf{el}(S)\})=O(\sigma\mathsf{e}_{\min})$.
@@ 4.2 Computing EBFs
Recall our algorithm for computing $\mathsf{MAW}(S)$ in Section 4.1. Let $u_{i}$ be a node that has (soft or hard) Weiner link with label $a\in\Sigma$ and has an out-edge starting with $b\in\Sigma$. Assuming $S[1]=\sharp$ and $S[|S|]=\$$, there are characters $a^{\prime},b^{\prime}\in\Sigma$ such that $a^{\prime}\neq a$, $b^{\prime}\neq b$, and $a^{\prime}u_{i},u_{i}b^{\prime}\in\mathsf{Substr}(S)$. Note also that $au_{i},u_{i}b\in\mathsf{Substr}(S)$. Thus, if $au_{i}b\in\mathsf{Substr}(S)$, then $au_{i}b\in\mathsf{EBF}(S)$. If the locus for $au_{i}$ is the node $\hat{u}$, in which case the Weiner link from $u_{i}~{}(=u_{1})$ to $\hat{u}$ is hard, then the EBF $au_{i}b$ is charged to the out-edge of $\hat{u}$ that starts with $b$. If the locus for $au_{i}$ is on the edge $(\hat{v},\hat{u})$, then the EBF $au_{i}b$ is charged to the soft-Weiner link from $u_{i}$ that labeled $a$. Since the number of hard Weiner links is $|V|-1$, the number of soft-Weiner links is $\mathsf{el}(S)-|V|+1$. Thus we have $|\mathsf{EBF}(S)|=\mathsf{er}(S)+\mathsf{el}(S)-|V|+1$ for any string $S$ with $S[1]=\sharp$ and $S[|S|]=\$$.
@@
By modifying our algorithm for computing MAWs for $S$ as is described in the poof for Lemma 5, we can obtain the following:
@@ 4.3 Computing length-bounded MAWs and EBFs
In our data structure for reporting $\mathsf{MAW}(S)^{\leq\ell}$, we precompute the largest integer $m$ with $|\mathsf{MAW}(S)^{\leq m}|\leq\mathsf{e}_{\min}$, and store the elements in $\mathsf{MAW}(S)^{\leq m}$ using Lemma 3, sorted in their lengths. Given a query length $\ell$, if $\ell\leq m$, then we can output the elements of $\mathsf{MAW}(S)^{\leq\ell}$ by reporting the stored MAWs in increasing order, in $O(|\mathsf{MAW}(S)^{\leq\ell}|+1)$ time. Otherwise (if $\ell>m$), then we process each edge $(\hat{v},\hat{u})$ in $\mathsf{LPT}^{+}(S)$ in increasing order of $|\mathsf{str}(\hat{u})|$. We need to access the internal nodes in the path from $\mathsf{slink}(\hat{v})$ to $\mathsf{slink}(\hat{u})$ topdown, in $O(1)$ time each. This can be done in $O(1)$ time per node by augmenting $\mathsf{LPT}^{+}(S)$ with the level-ancestor (LA) data structure [8, 6]. We stop processing the edge $(\hat{v},\hat{u})$ as soon as we find the shallowest ancestor $u_{i}$ of $u=\mathsf{slink}(\hat{u})$ such that $|\mathsf{str}(u_{i})|>\ell-2$. Since $|\mathsf{MAW}(S)^{\leq\ell}|>\mathsf{e}_{\min}$ for any $\ell<m$, this algorithm outputs $\mathsf{MAW}(S)^{\leq\ell}$ in $O(|\mathsf{MAW}(S)^{\leq\ell}|+1)$ time.
Computing $\mathsf{MAW}(S)^{\geq\ell}$ is symmetric, except that the LA data structure is not needed: We can check the nodes between $\mathsf{slink}(\hat{u})$ and $\mathsf{slink}(\hat{v})$ in deceasing order of their lengths, in $O(1)$ time each, by simply climbing up the path. ∎
@@ 5 Computing MRWs
For any $w\in\bigcup_{k\geq 1}\mathsf{MRW}_{k}(S)$, the locus of $w$ is the position on some edge $(v_{1},v_{2})$ exactly one character down from $v_{1}$ since $\mathsf{occ}_{S}(w[1..|w|-1])>\mathsf{occ}_{S}(w)$. For the sake of contradiction, we assume that the locus of $w_{1}=a_{1}u_{1}b\in\bigcup_{k\geq 1}\mathsf{MRW}_{k}(S)$ is the same as that of another $w_{2}=a_{2}u_{2}b\in\bigcup_{k\geq 1}\mathsf{MRW}_{k}(S)$ where $a_{1},a_{2},b\in\Sigma$ and $u_{1},u_{2}\in\Sigma^{\star}$. We further assume $|w_{1}|\leq|w_{2}|$ w.l.o.g. Since $a_{1}u_{1}$ and $a_{2}u_{2}$ reach the same CDAWG node, $a_{1}u_{1}$ is a suffix of $a_{2}u_{2}$. Since $w_{1}\neq w_{2}$, $|w_{1}|<|w_{2}|$ holds, and thus, $w_{1}$ is a proper suffix of $w_{2}$. This contradicts the minimality of MRW $w_{2}$. Hence, the number of MRWs in $\bigcup_{k\geq 1}\mathsf{MRW}_{k}(S)$ is never larger than $|\mathsf{E}|$. The above argument holds even if the reversal of $S$ is considered. Therefore, the upper bound is $\min\{|\mathsf{E}|,|\mathsf{E}^{R}|\}=\mathsf{e}_{\min}$. ∎

### v2

@@ Abstract
A string $w$ is said to be a minimal absent word (MAW) for a string $S$ if $w$ does not occur in $S$ and any proper substring of $w$ occurs in $S$. We focus on non-trivial MAWs which are of length at least 2. Finding such non-trivial MAWs for a given string is motivated by applications in bioinformatics and data compression. Fujishige et al. [TCS 2023] proposed a data structure of size $\Theta(n)$ that can output the set $\mathsf{MAW}(S)$ of all MAWs for a given string $S$ of length $n$ in $O(n+|\mathsf{MAW}(S)|)$ time, based on the directed acyclic word graph (DAWG). In this paper, we present a more space efficient data structure based on the compact DAWG (CDAWG), which can output $\mathsf{MAW}(S)$ in $O(|\mathsf{MAW}(S)|)$ time with $O(\mathsf{e}_{\min})$ space, where $\mathsf{e}_{\min}$ denotes the minimum of the sizes of the CDAWGs for $S$ and for its reversal $S^{R}$. For any strings of length $n$, it holds that $\mathsf{e}_{\min}<2n$, and for highly repetitive strings $\mathsf{e}_{\min}$ can be sublinear (up to logarithmic) in $n$. We also show that MAWs and their generalization minimal rare words have close relationships with extended bispecial factors, via the CDAWG.
@@ 1 Introduction
A string $w$ is said to be a minimal absent word (MAW) for a string $S$ if $w$ does not occur in $S$ and any proper substring of $w$ occurs in $S$. Throughout this paper, we focus on MAWs which are of form $aub$, where $a,b$ are characters and $u$ is a (possibly empty) string. Finding such MAWs for a given string is motivated by applications including bioinformatics [1, 12, 23, 19] and data compression [14, 15, 3].
@@
In this paper, we present a more space efficient data structure based on the compact DAWG (CDAWG[10], which takes $O(\mathsf{e}_{\min})$ space and can output $\mathsf{MAW}(S)$ in $O(|\mathsf{MAW}(S)|)$ time, where $\mathsf{e}_{\min}=\min\{\mathsf{el}(S),\mathsf{er}(S)\}$, with $\mathsf{el}(S)$ and $\mathsf{er}(S)$ being the numbers of edges of the CDAWGs for $S$ and for its reversal $S^{R}$, respectively. For any string $S$ of length $n$ it holds that the number of edges of the CDAWG for $S$ is less than $2n$ and it can be sublinear (up to logarithmic) in $n$ for highly repetitive strings [25, 24, 5].
@@ 2.1 Basic Notations
For a set $\mathsf{S}$ of strings, let $|\mathsf{S}|$ denote the number of strings in $\mathsf{S}$.
@@ 2.2 MAW, MUS, MRW, and EBF
A string $w=aub\in\mathsf{Substr}(S)$ with $a,b\in\Sigma$ and $u\in\Sigma^{*}$ is called an extended bispecial factor (EBF) for string $S$ if $\mathsf{occ}_{S}(a^{\prime}u)\geq 1$ and $\mathsf{occ}_{S}(ub^{\prime})\geq 1$ for some characters $a^{\prime}\neq a$ and $b^{\prime}\neq b$. Let $\mathsf{EBF}(S)$ denote the set of EBFs for string $S$.
@@ 4.1 Computing MAWs
Let $u$ be any node in $\mathsf{CDAWG}(S)$ and suppose it has $k$ in-coming edges. The above process of adding secondary edges to $\mathsf{LPT}(S)$ can be regarded as splitting $u$ into $k+1$ nodes $u^{\prime},\ell_{1},\ldots,\ell_{k}$, such that $|\mathsf{str}(u^{\prime})|>|\mathsf{str}(\ell_{1})|>\cdots>|\mathsf{str}(\ell_{k})|$. We define the suffix links of these nodes such that $\mathsf{slink}(u^{\prime})=\ell_{1}$, $\mathsf{slink}(\ell_{i})=\ell_{i+1}$ for $1\leq i<k$, and $\mathsf{slink}(\ell_{k})=v^{\prime}$ such that $v^{\prime}$ is the node in $\mathsf{LPT}^{+}(S)$ that corresponds to the destination node $v=\mathsf{slink}(u)$ of the suffix link of $u$ in $\mathsf{CDAWG}(S)$. We call these copied nodes $\ell_{1},\ldots,\ell_{k}$ as gray nodes, and the other node $u^{\prime}$ as a white node. For convenience, we assume that the source is a white node. See Fig. 3 for illustration.
@@
For $i=1$, node $u_{1}~{}(=u)$ has a hard Weiner link to $\hat{u}$. Let $a$ be the character label of this hard Weiner link. Let $\mathsf{child}(\hat{u})$ and $\mathsf{child}(\mathsf{slink}(\hat{u}))=\mathsf{child}(u_{1})$ be the sorted lists of the first characters of the out-edge labels from $\hat{u}$ and from $u_{1}$. Note that $\mathsf{child}(\hat{u})\subseteq\mathsf{child}(u_{1})$. Then, it follows from the definition of MAWs and Lemma 1 that $au_{1}b\in\mathsf{MAW}(S)$ iff $b\in\mathsf{child}(u_{1})\setminus\mathsf{child}(\hat{u})$.
For each $i=2,\ldots,m$, observe that node $u_{i}$ has a (virtual) soft Weiner link with the same character label $a$, and leading to the same node $\hat{u}$. We emphasize that this soft Weiner link is only virtual and is not stored in our $O(\mathsf{er}(S))$-space data structure. We have the following lemma:
@@
Let $c$ be the first character of the label of the edge $(u_{i+1},u_{i})$. Let $\hat{u}_{i}$ denote the (virtual) implicit node on the edge $(\hat{v},\hat{u})$, such that $\mathsf{slink}(\hat{u}_{i})=u_{i}$. Note that $c$ is the unique character that immediately follows the locus of $\hat{u}_{i}$ in the string label of edge $(\hat{v},\hat{u})$. Since $u_{i}$ is branching but $\hat{u}_{i}$ is non-branching, the set $\mathsf{child}(u_{i})\setminus\mathsf{child}(\hat{u}_{i})=\mathsf{child}(u_{i})\setminus\{c\}$ is not empty. Thus, for any character $b\in\mathsf{child}(u_{i})\setminus\{c\}$, we have $au_{i}b\in\mathsf{MAW}(S)$. ∎
@@
Let us analyze the time complexity and the data structure space usage. In the case with $i\geq 2$, by Lemma 4, accessing each node $u_{i}$ can be charged to any one of the MAWs reported with $u_{i}$. The time required for $u_{i}$ is thus linear in the number of reported MAWs $au_{i}b$. In the case with $i=1$, we compare the sorted lists $\mathsf{child}(\hat{u})$ and $\mathsf{child}(u_{1})$. Each successful character comparison with a character $b\in\mathsf{child}(u_{1})\setminus\mathsf{child}(\hat{u})$ reports a MAW $au_{1}b$. On the other hand, each unsuccessful character comparison with a character $b\in\mathsf{child}(u_{1})\cap\mathsf{child}(\hat{u})$ can be charged to the corresponding out-edge of $\hat{u}$ whose string label begins with $b$. Thus, we can output all MAWs in $O(\mathsf{er}(S)+|\mathsf{MAW}(S)|)$ time. Further, when $|\mathsf{MAW}(S)|\leq\mathsf{er}(S)$, then we can store the set $\mathsf{MAW}(S)$ with $O(|\mathsf{MAW}(S)|)\subseteq O(\mathsf{er}(S))$ space and output each MAW in $O(1)$ time, using Lemma 3. This leads to our $O(|\mathsf{MAW}(S)|)$-time $O(\mathsf{er}(S))$-space data structure.
@@
For each node $u_{i}$ of $\mathsf{LPT}^{+}(S)$, the number of distinct characters $a$ such that $au_{i}b\in\mathsf{MAW}(S)$ is at most the number of (hard or soft) Weiner links from $u_{i}$, which is bounded by $\sigma$. The number of distinct characters $b$ such that $au_{i}b\in\mathsf{MAW}(S)$ does not exceed the number of out-edges from $u_{i}$. By summing them up for all nodes $u_{i}$ of $\mathsf{LPT}^{+}(S)$, we obtain $|\mathsf{MAW}(S)|=O(\sigma\cdot\mathsf{er}(S))$. Since the same argument holds for the reversed string $S^{R}$ and $\mathsf{LPT}^{+}(S^{R})$, we have $|\mathsf{MAW}(S)|=O(\sigma\cdot\min\{\mathsf{er}(S),\mathsf{el}(S)\})=O(\sigma\mathsf{e}_{\min})$.
@@ 4.2 Computing EBFs
Recall our algorithm for computing $\mathsf{MAW}(S)$ in Section 4.1. Let $u_{i}$ be a node that has a (soft or hard) Weiner link with label $a\in\Sigma$ and has an out-edge starting with $b\in\Sigma$. Assuming $S[1]=\sharp$ and $S[|S|]=\$$, there are characters $a^{\prime},b^{\prime}\in\Sigma$ such that $a^{\prime}\neq a$, $b^{\prime}\neq b$, and $a^{\prime}u_{i},u_{i}b^{\prime}\in\mathsf{Substr}(S)$. Note also that $au_{i},u_{i}b\in\mathsf{Substr}(S)$. Thus, if $au_{i}b\in\mathsf{Substr}(S)$, then $au_{i}b\in\mathsf{EBF}(S)$. If the locus for $au_{i}$ is the node $\hat{u}$, in which case the Weiner link from $u_{i}~{}(=u_{1})$ to $\hat{u}$ is hard, then the EBF $au_{i}b$ is charged to the out-edge of $\hat{u}$ that starts with $b$. If the locus for $au_{i}$ is on the edge $(\hat{v},\hat{u})$, then the EBF $au_{i}b$ is charged to the soft-Weiner link from $u_{i}$ that labeled $a$. Since the number of hard Weiner links is $|V|-1$, the number of soft-Weiner links is $\mathsf{el}(S)-|V|+1$. Thus we have $|\mathsf{EBF}(S)|=\mathsf{er}(S)+\mathsf{el}(S)-|V|+1$ for any string $S$ with $S[1]=\sharp$ and $S[|S|]=\$$.
@@
By modifying our algorithm for computing MAWs for $S$ as is described in the proof for Lemma 5, we can obtain the following:
@@ 4.3 Computing length-bounded MAWs and EBFs
In our data structure for reporting $\mathsf{MAW}(S)^{\leq\ell}$, we precompute the largest integer $m$ with $|\mathsf{MAW}(S)^{\leq m}|\leq\mathsf{e}_{\min}$, and store the elements in $\mathsf{MAW}(S)^{\leq m}$ using Lemma 3, sorted by their lengths. Given a query length $\ell$, if $\ell\leq m$, then we can output the elements of $\mathsf{MAW}(S)^{\leq\ell}$ by reporting the stored MAWs in increasing order, in $O(|\mathsf{MAW}(S)^{\leq\ell}|+1)$ time. Otherwise (if $\ell>m$), then we process each edge $(\hat{v},\hat{u})$ in $\mathsf{LPT}^{+}(S)$ in increasing order of $|\mathsf{str}(\hat{u})|$. We need to access the internal nodes in the path from $\mathsf{slink}(\hat{v})$ to $\mathsf{slink}(\hat{u})$ topdown, in $O(1)$ time each. This can be done in $O(1)$ time per node by augmenting $\mathsf{LPT}^{+}(S)$ with the level-ancestor (LA) data structure [8, 6]. We stop processing the edge $(\hat{v},\hat{u})$ as soon as we find the shallowest ancestor $u_{i}$ of $u=\mathsf{slink}(\hat{u})$ such that $|\mathsf{str}(u_{i})|>\ell-2$. Since $|\mathsf{MAW}(S)^{\leq\ell}|>\mathsf{e}_{\min}$ for any $\ell<m$, this algorithm outputs $\mathsf{MAW}(S)^{\leq\ell}$ in $O(|\mathsf{MAW}(S)^{\leq\ell}|+1)$ time.
Computing $\mathsf{MAW}(S)^{\geq\ell}$ is symmetric, except that the LA data structure is not needed: We can check the nodes between $\mathsf{slink}(\hat{u})$ and $\mathsf{slink}(\hat{v})$ in decreasing order of their lengths, in $O(1)$ time each, by simply climbing up the path. ∎
@@ 5 Computing MRWs
For any $w\in\bigcup_{k\geq 1}\mathsf{MRW}_{k}(S)$, the locus of $w$ is the position on some edge $(v_{1},v_{2})$ exactly one character down from $v_{1}$ since $\mathsf{occ}_{S}(w[1..|w|-1])>\mathsf{occ}_{S}(w)$. For the sake of contradiction, we assume that the locus of $w_{1}=a_{1}u_{1}b\in\bigcup_{k\geq 1}\mathsf{MRW}_{k}(S)$ is the same as that of another $w_{2}=a_{2}u_{2}b\in\bigcup_{k\geq 1}\mathsf{MRW}_{k}(S)$ where $a_{1},a_{2},b\in\Sigma$ and $u_{1},u_{2}\in\Sigma^{\star}$. We further assume $|w_{1}|\leq|w_{2}|$ w.l.o.g. Since $a_{1}u_{1}$ and $a_{2}u_{2}$ reach the same CDAWG node, $a_{1}u_{1}$ is a suffix of $a_{2}u_{2}$. Since $w_{1}\neq w_{2}$, $|w_{1}|<|w_{2}|$ holds, and thus, $w_{1}$ is a proper suffix of $w_{2}$. This contradicts that $\mathsf{occ}_{S}(w_{2})=\mathsf{occ}_{S}(w_{1})$ and $w_{2}$ is an MRW. Hence, the number of MRWs in $\bigcup_{k\geq 1}\mathsf{MRW}_{k}(S)$ is never larger than $|\mathsf{E}|$. The above argument holds even if the reversal of $S$ is considered. Therefore, the upper bound is $\min\{|\mathsf{E}|,|\mathsf{E}^{R}|\}=\mathsf{e}_{\min}$. ∎